# Collision-Free Trajectory Optimization in Cluttered Environments with Sums-of-Squares Programming

### v1

@@ IV-C Collision-Free Trajectory Optimization with Implicit Safety Constraints
With the geometrical interpretation of the containment relationship in (2), we define a uniformly scaling factor $\alpha_{\tau}$ for each waypoint $\boldsymbol{q}_{\tau}$ in $\mathcal{X}$, and reformulate the collision-free trajectory optimization problem (III) as:
@@
In (IV-C), motion safety considering geometry of the robot is ensured by enforcing $\alpha_{\tau}$ to be no greater than one over the entire optimization horizon, where $\alpha_{\tau}(\boldsymbol{q}_{\tau})$ is the optimum of the proposed minimum scaling problem (IV-B2), which is implicitly dependent on the robot configuration $\boldsymbol{q}$ at time stamp $\tau$. In this section, to deal with the implicit safety constraints, we further derive the gradient of the scaling factor with respect to the robot configuration $\partial{\alpha}/\partial{\boldsymbol{q}}\in\mathbb{R}^{n}$ analytically. This facilitates the seamless incorporation of the proposed optimization problem (7) into gradient-based trajectory optimization algorithms, thereby enabling the effective solving of the nonlinear program (IV-C) with the proposed implicit safety constraints.
In (7), $\alpha$ is simply the objective function, and in this case, $\partial{\alpha}/\partial{\boldsymbol{q}}$ can be effectively extracted by calculating the partial derivative of the Lagrangian function $\mathcal{L}$ with respect the robot configuration $\boldsymbol{q}$ at the primal-dual optimum of the SDP (7). Furthermore, since the rotation matrix $R(\boldsymbol{q})$ and translation vector $p(\boldsymbol{q})$ only appear in the coefficients of $f^{\mathcal{Q}}$, it is only necessary to consider the Lagrange multipliers corresponding to the constraints in (7a) to compute the gradient. In this sense, the gradient vector can be represented as:
$\frac{\partial{\alpha}}{\partial\boldsymbol{q}}$ $=\frac{\partial\mathcal{L}}{\partial\boldsymbol{q}}=\sum_{i,l}\lambda_{i,l}\frac{\partial{\big{(}{\rm coef}(f^{\mathcal{Q}_{\tau}}_{i})_{l}}\big{)}}{\partial\boldsymbol{q}},$ (13)
@@
Remark 3: Under generic optimality conditions, optimal solutions of (7) and (IV-B2) are identical for all $k$ that is sufficiently large [18]. This situation can be further improved if all the polynomial functions involved are SOS-convex [16], under which condition the SOS relaxation at the lowest order is exact.
Subsequentlly, we employ the gradient-based algorithm Augmented Lagrangian Iterative Linear Quadratic Regulator (AL-iLQR) proposed in [19]. In this approach, the constraints can be augmented into the objective function and solved iteratively using the iLQR algorithm. Specifically, we calculate the value of $\alpha$ as well as its gradient with respect to the robot configuration $\boldsymbol{q}$ based on the trajectory output from each round of forward pass of AL-iLQR. We then integrate this information in the execution of the backward pass of AL-iLQR to calculate the perturbed action, enforcing the robot to stay inside the free regions $\mathcal{Q}$. Notably, the process of solving the proposed SDP and extracting gradients along the trajectory is significantly accelerated through precomputation and parallelization.
To this end, the proposed collision-free trajectory optimization method considering specific robot geometries is summarized in Algorithm 1.
@@ 1 Function WaypointsAllocation($\mathcal{Q},\boldsymbol{q}_{s},\boldsymbol{q}_{g},T$):
3
4Initialization: $\mathcal{X}^{*}\leftarrow\mathcal{X}_{0}$ and $U^{*}\leftarrow U_{0}$;$ctol$leftarrow Tolerance of constraints violation in AL-iLQR;c $\leftarrow$ Constraints violation of current trajectory; Index $\leftarrow$ WaypointsAllocation($\mathcal{Q},\boldsymbol{q}_{s},\boldsymbol{q}_{g},T$); while $\texttt{c}\geq ctol$ do
@@
8      $\delta u\leftarrow$ ALiLQR-BACKWARDPASS($\mathcal{X}^{*}$, $U^{*}$, $\alpha$, $\frac{\partial\alpha}{\partial\boldsymbol{q}}$); $\mathcal{X}^{*},U^{*},\texttt{c}\leftarrow$ALiLQR-FORWARDPASS($\mathcal{X}^{*},\,U^{*},\,\delta u$)
@@ V Results
In this section, we conduct a thorough evaluation of our proposed trajectory optimization approach through simulations and real-world experiments. We deploy the algorithm on an Intel i5-13400F processor, and the proposed SDP in (7) is implemented and solved parallelly using MOSEK [20] with precomputation of the pertinent parameters. In simulations, we initially validate our algorithm on robots modeled with various geometries represented as semialgebraic sets, including a polytope, a cylinder, an ellipsoid, and an elliptical cone. We analyze typical scenarios where waypoints fail to converge into the free regions and discuss how we address these issues with specific ingredients. Subsequently, we demonstrate the efficacy and robustness of our algorithm for collision-free trajectory optimization in three challenging maze scenarios with different degrees of complexity. Finally, we apply the trajectory optimization pipeline on the real robot platform Unitree Go1, which showcases the adaptability and practicality of our method for safely navigating through a cluttered indoor environment. The entire process is implemented on an intel NUC11 with i5-7260U processor.
@@ V-A1 Algorithm Validation for Various Robot Geometries
In this scenario, we validate the effectiveness of our proposed safety constraints on robots with different geometries. We initialize the trajectories at random configurations and command the robots to stop within the safe region by enforcing the terminal safety constraint, i.e., $\alpha_{T}(\boldsymbol{q}_{T})\leq 1$. As shown in Fig. 3, trajectories for robots with all geometries have successfully been “attracted” into the designated region with the help of the proposed minimal scaling problem (7) as well as its gradients information (13).
The semialgebraic set descriptions for each involved robot geometry are listed in Table II, which is defined in their body frames as shown in Fig. 3. Note that the polynomial functions defining the geometries are all SOS-convex. We achieve 12$\pm$5 ms in average for solving the proposed SDP at the minimum relaxation order since the polynomial functions defining the geometries are all SOS-convex, and the time spent on formulating the SDP and calculating the gradient can be neglected as the explicit expressions have been pre-computed.
Several ingredients are added in our method to ensure the convergence of (IV-C) with our derived safety constraints. For the SDP (7), the Lagrange multiplier $\lambda_{i,l}$ is zero unless the corresponding constraint is active at the primal-dual optimum, as elaborated in the complementary slackness of KKT optimality conditions [21]. Typically, as indicated in Fig. 2, only the constraints corresponding to the facet that restricts further decrease of $\alpha$ is active at the optimum of the proposed minimum scaling problem. Thus, we use the information of the most difficult-to-satisfy facet for the gradients computation (13) to push the robot towards a safer configuration. This mechanism fails to efficiently utilize all the information from the violated faces to collectively push the robot into the safe region, and could potentially lead to unexpected local optima during line search process, due to the discontinuity from facet switching. To address this problem, we extract the gradients of all violated facets and synthesize them into a more representative gradient by continuously relaxing the constraints of the corresponding facets in (7a) and re-solving the SDP. Another observation from the results is that the gradient scale for translational motion is larger than that for rotational motion. This results in negligible changes in the rotation angle after line search if both rotational and translational motions are searched simultaneously, even if a better trajectory exists with larger changes in the rotation angle. In this sense, we adopt a two-layer line search scheme for translation and rotation separately to avoid trapping into the local optimum and ensure continuous convergence.
@@
TABLE II: Geometry Descriptions of the sets $\mathcal{W}_{\mathcal{B}}$ for different Geometries
Geometries $\mathcal{W}_{\mathcal{B}}$
Polytope $\{\boldsymbol{x}=(x,y,z):A\boldsymbol{x}\leq\boldsymbol{b}\}$
Ellipsoid $\{\boldsymbol{x}=(x,y,z):\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}\leq 1\}$
Elliptical Cylinder $\{\boldsymbol{x}=(x,y,z):\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}\leq 1,z\geq-\frac{h}{2},z\leq\frac{h}{2}\}$
Elliptical Cone $\{\boldsymbol{x}=(x,y,z):\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}\leq\frac{z^{2}}{c^{2}},z\geq-\frac{h}{2},z\leq\frac{h}{2}\}$
Figure 4: Performance of our proposed trajectory optimization framework in maze traversing tasks. (a-c) Three maze scenarios with different degrees of complexity. The start and goal positions are denoted by red and green nodes, respectively. (d-f) Visualization of the collision-free trajectories generated by the proposed method in all the three maze scenarios with key frames highlighted.
@@ V-A2 Traversing in Mazes
To evaluate the performance of our proposed trajectory optimization framework in challenging scenarios, we have crafted three distinct mazes of varying dimensions and degrees of complexity, namely the large, medium, and small mazes, respectively, as shown in Fig. 4(a)-4(c). To demonstrate the significance of the two proposed ingredients, we conduct two sets of ablation experiments: one with our method excluding gradient synthesis (Ours w/o GS), and the other with our method excluding the two-layer line search (Ours w/o TLS). We record the optimization horizons in the three maze scenarios, as well as the number of iterations before the AL-iLQR cost ceases to decrease for the three different experimental settings, and the final count of waypoints outside the safe region. Observed from Table III, it is evident that as the complexity of the scene increases, the absence of these two ingredients makes it more susceptible to local optimum, which leads to an inability to further reduce the cost associated with safety constraints, thereby resulting in unsafe situations. Remarkably, our method successfully optimizes collision-free trajectories in all the three scenarios with the smallest number of iterations. As depicted in the Fig. 4(d)-4(f), our optimized trajectories enable the robot to navigate through diverse complex maze environments by adopting different degrees of posture adjustments based on its shape, thereby demonstrating the effectiveness of our collision-free trajectory optimization method.
TABLE III: Comparisons between our method with or without two key ingredients in maze scenarios
Maze Waypoints Method Iterations Collisions
Large 143 Ours w/o GS 15 9
Ours w/o TLS 18 13
Ours 16 0
Medium 70 Ours w/o GS 17 7
Ours w/o TLS 13 9
Ours 11 0
Small 16 Ours w/o GS 6 2
Ours w/o TLS 7 0
Ours 5 0
@@ V-B Real-World Experiment
In this section, we demonstrate the practicality and robustness of our proposed method by deploying it on a real robot platform Unitree Go1. In the experiment, the quadruped is commanded to traverse through a cluttered indoor area of $2\,\textup{m}\times 6\,\textup{m}$ with several randomly placed obstacles and a hole to simulate the height-constrained situation.
We use the onboard Robosense 16-line spinning Lidar to perceive point cloud data for off-line space decomposition and real-time align-based localization. We incorporate additional physical constraints of the quadruped in the trajectory optimization process and implement the kinematic feedforward and feedback controllers introduced in [22] for trajectory tracking. The computed body twist commands are subsequently fed into the embedded speed controller to actuate each joint on the robot. The robot, along with the onboard sensor, is approximated as a polytope. The proposed optimization problem (7) is solved in around 15$\pm$3 ms. As depicted in Fig. 5, the quadruped adjusts its pose to adapt to the obstacle configurations, and due to the tightly fitted robot geometry, the quadruped lowers its pose to successfully traverse the height-constrained hole in a non-conservative manner (shown in Fig. 5(g)-5(i)). Overall, our proposed framework exhibits remarkable capacity for generating safe and flexible maneuvers in the presence of complex obstacle configurations.

### v2

@@ IV-C Collision-Free Trajectory Optimization with Implicit Safety Constraints
With the geometrical interpretation of the containment relationship in (2), we define a uniformly scaling factor $\alpha_{\tau}$ at each waypoint $\boldsymbol{q}_{\tau}$ in $\mathcal{X}$, and reformulate the collision-free trajectory optimization problem (III) as:
@@
In (IV-C), motion safety considering geometry of the robot is ensured by enforcing $\alpha_{\tau}$ to be no greater than one over the entire optimization horizon, where $\alpha_{\tau}(\boldsymbol{q}_{\tau})$ is the optimum of the proposed minimum scaling problem (IV-B2), which is implicitly dependent on the robot configuration $\boldsymbol{q}$ at time stamp $\tau$.
In this section, to handle the implicit safety constraints, we calculate a direction for updating the robot configuration $q$ to decrease the value function $\nu(\boldsymbol{q})$ of (7), i.e., the minimum scaling factor $\alpha$ in (7) with the given robot configuration $q$. In practical implementation, we use $-\partial{\mathcal{L}}/\partial{\boldsymbol{q}}$ at the primal-dual solution of (7) to update $\boldsymbol{q}$, which performs greatly in computation and has also been widely exploited for trajectory optimization in robotics [9, 23]. Indeed, one may show that under mild conditions, the gradient $\partial{\mathcal{L}}/\partial{\boldsymbol{q}}$ belongs to an upper approximation for the Clarke subgradient of $\nu(\boldsymbol{q})$; see [24, Corollary 1, page 242]. Furthermore, if we additionally suppose $\nu(\boldsymbol{q})$ is differentiable (this is the case when (7) has a unique primal-dual solution and Slater’s condition holds), then $\nabla\nu=\partial{\mathcal{L}}/\partial{\boldsymbol{q}}$, by [24, Corollary 2, page 242].
In (7), since the rotation matrix $R(\boldsymbol{q})$ and translation vector $p(\boldsymbol{q})$ appear in the coefficients ${\rm coef}(f^{\mathcal{Q}_{\tau}}_{i}(\alpha))_{l}$, it is only necessary to consider the terms in (7a) when computing the gradient $\partial\mathcal{L}/\partial\boldsymbol{q}$. Therefore, the gradient vector can be represented as:
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{q}}=\sum_{i,l}\lambda_{i,l}\frac{\partial{\big{(}{\rm coef}(f^{\mathcal{Q}_{\tau}}_{i})_{l}}\big{)}}{\partial\boldsymbol{q}},$ (13)
@@
Subsequently, we employ the gradient-based algorithm Augmented Lagrangian Iterative Linear Quadratic Regulator (AL-iLQR) proposed in [5]. In this approach, the constraints can be augmented into the objective function and solved iteratively using the iLQR algorithm. Specifically, we calculate the value of $\alpha$ as well as the gradient information based on the trajectory $\boldsymbol{q}$ output from each round of forward pass of AL-iLQR. We then integrate this information in the execution of the backward pass of AL-iLQR to calculate the perturbed action, forcing the robot to stay inside the free regions $\mathcal{Q}$. To this end, the proposed collision-free trajectory optimization method considering specific robot geometries is summarized in Algorithm 1.
Remark 3: In the proposed bi-level optimization process for solving (IV-C), safety is ensured by enforcing the minimal scaling factor $\alpha$ to be less than 1 along the trajectory and using this as a termination criterion for the AL-iLQR to converge.
@@ 1 Function WaypointsAllocation($\mathcal{Q},\boldsymbol{q}_{s},\boldsymbol{q}_{g},T$):
2       $\mathcal{Q}\leftarrow$Approximate $\mathcal{F}$ with sequence of free regions;
3       $\mathcal{V}\leftarrow$ Assign each region $\mathcal{Q}_{i}$ in $\mathcal{Q}$ to a vertex $v_{i}$;
4       $\mathcal{E}\leftarrow$ Connect ($v_{i}$,$v_{j}$) in $\mathcal{V}$ if they are overlapped;
5       $\mathbb{P}^{*}_{v}\leftarrow$ Search for free regions sequence connecting $\boldsymbol{q}_{s}$ and $\boldsymbol{q}_{g}$ with shortest path length;
6       $\mathbb{P}^{*}_{e}\leftarrow$ Extract the edge sequence concatenating vertexes in $\mathbb{P}^{*}_{v}$;
@@ 4       $\mathcal{E}\leftarrow$ Connect ($v_{i}$,$v_{j}$) in $\mathcal{V}$ if they are overlapped;
[…]
9       return Index
10
11Initialization:
12 $\mathcal{X}^{*}\leftarrow\mathcal{X}_{0}$ and $U^{*}\leftarrow U_{0}$;
[…]
14c $\leftarrow$ Constraints violation of current trajectory;
15 Index $\leftarrow$ WaypointsAllocation($\mathcal{Q},\boldsymbol{q}_{s},\boldsymbol{q}_{g},T$);
16 while $\texttt{c}\geq ctol$ do
@@ 15 Index $\leftarrow$ WaypointsAllocation($\mathcal{Q},\boldsymbol{q}_{s},\boldsymbol{q}_{g},T$);
20      $\delta u\leftarrow$ ALiLQR-BACKWARDPASS($\mathcal{X}^{*}$, $U^{*}$, $\alpha$, $\frac{\partial\alpha}{\partial\boldsymbol{q}}$);
21       $\mathcal{X}^{*},U^{*},\texttt{c}\leftarrow$ALiLQR-FORWARDPASS($\mathcal{X}^{*},\,U^{*},\,\delta u$)
@@ V Results
In this section, we thoroughly evaluate our proposed trajectory optimization approach through simulations and real-world experiments. For simulations, we run the algorithm on an Intel i5-13400F processor, and the proposed SDP in (7) is implemented and solved parallelly using MOSEK [25] with precomputation of the pertinent parameters. Initially, we validate our algorithm on robots modeled with various geometries represented as semialgebraic sets. We then demonstrate the efficacy and robustness of our algorithm for collision-free trajectory optimization in three challenging maze scenarios with varying levels of complexity. Also, we conduct comparative experiments to further showcase the safety and computational advantages of our method in complex and obstacle-dense environments. Finally, we apply the trajectory optimization pipeline on the real robot platform Unitree Go1, with the entire process implemented on an intel NUC11 with i5-7260U processor, highlighting the adaptability and practicality of our method for safely navigating through a cluttered indoor environment.
@@ V-A1 Algorithm Validation for Various Robot Geometries
In this scenario, we validate the effectiveness of our proposed safety constraints on robots with different geometries represented as semi-algebraic sets, including a polytope, a cylinder, an ellipsoid, and an elliptical cone. We initialize the trajectories at random configurations and command the robots to stop within the safe region by enforcing the terminal safety constraint, i.e., $\alpha_{T}(\boldsymbol{q}_{T})\leq 1$. As shown in Fig. 3, trajectories for robots with all geometries have successfully been “attracted” into the designated region utilizing the proposed bi-level optimization framework with the scaling-based safety constraints. We achieve 10 $\pm$ 5 ms in average for solving the proposed SDP at the minimum relaxation order since the polynomial functions defining the geometries are all SOS-convex, and the time spent on formulating the SDP and calculating the gradient can be neglected as the explicit expressions have been pre-computed.
@@
Figure 4: Performance of our proposed trajectory optimization framework in maze traversing tasks. (a-c) Three maze scenarios with different degrees of complexity. The start and goal positions are denoted by red and green nodes, respectively. (d-f) Visualization of the collision-free trajectories generated in the graph of free regions $\mathcal{G}$ with keyframes highlighted.
(a) Kinodynamic RRT*[2]
(b) TrajOpt[14]
(c) CBF-Polytopes[7]
(d) Safe-Corridor[10]
(e) Ours
Figure 5: Comparsion results in a 5 m×$\times$10 m cluttered environment. A 0.2 m $\times$ 0.4 m polytopic robot is commanded to move from initial (red node) to goal configuration (green node). (a) Visualization of the search process of Kinodynamic RRT*. It fails to find a feasible trajectory through the narrow passages (highlighted) within tractable time (100 s). (b-c) Optimized trajectories of TrajOpt and CBF-Polytopes. They fail to push the robot away from all the obstacles simultaneously within the optimization horizon, and the collision areas are highlighted. (d-e) Visualization of the generated free regions and the planned trajectories of Safe-Corridor and our proposed method. Both approaches generate safe trajectories by decomposing free space. However, Safe-Corridor searches for paths and extracts free regions based on obstacle information that is dilated by the robot’s maximum side size. This method fails to find the nearest traversable path. Additionally, as highlighted by the gaps between the generated free regions and the obstacles in (d), the free regions do not cover enough of the available free space, resulting in a conservative trajectory.
@@ V-A2 Traversing in Mazes
To evaluate the performance of our proposed trajectory optimization framework in challenging scenarios, we have crafted three distinct mazes of varying dimensions and degrees of complexity, namely the small, medium, and large mazes, respectively, as shown in Fig. 4(a)-4(c). Remarkably, our method successfully optimizes collision-free trajectories in all three scenarios with fast convergence, taking 0.48 s, 2.5 s, and 4.2 s, respectively. As depicted in Fig. 4(d)-4(f), our optimized trajectories enable the robot to navigate through diverse complex maze environments by adopting different degrees of posture adjustments based on its shape, thereby demonstrating the effectiveness of our collision-free trajectory optimization method.
TABLE II: Comparisons between our method and benchmark methods in a cluttered maze environment
Methods Trajectory Length (m) Computation Time (s) Safety Non-conservative
 Map Processing Optimization
Kinodynamic RRT*[2] - - 100+ $\times$ $\times$
TrajOpt[14] 13.95 0.013 6.925 $\times$
CBF-Polytopes[7] 13.45 0.013 33.403 $\times$
Safe-Corridor[10] 17.97 0.018 0.256 $\times$
Ours 13.70 0.031 0.561
V-A3 Comparisons
We benchmark our proposed framework against Kinodynamic RRT*[2], TrajOpt[14], CBF-Polytopes[7], and Safe-Corridor[10]. The objective is to plan motions for a robot with polytopic geometry from the initial state to the goal state in a cluttered environment with complex obstacle layouts. Each of these methods differs from our approach in how it handles the environmental (obstacles) information, ensures safety, and solves the problem. For TrajOpt and CBF-Polytopes, we use the jump points search (JPS) algorithm to generate path references with no obstacle dilation, while in Safe-Corridor, obstacles are dilated based on the robot’s maximum side length in searching the reference path. Subsequently, it formulates a minimum snap QP problem and confines waypoints on the trajectory within the free regions extracted along the path. We compare computation time, trajectory length, and safety among the methods, with results recorded in Table II. The trajectories generated by each method are shown in Fig. 5.
As indicated in the Fig. 5(a)-5(c), Kinodynamic RRT*, TrajOpt, and CBF-Polytopes all failed to optimize safe trajectories. Kinodynamic RRT*, being a sampling-based method, struggles to generate valid paths due to the limited number of samples that fall in constrained spaces, reducing the probability of randomly sampling feasible paths through narrow passages. TrajOpt pushes the robot toward safe areas using locally calculated signed distances from the obstacle information, which can easily lead to getting stuck in local optima when facing complex obstacle layouts[3]. In the implementation of CBF-Polytopes, deriving CBF constraints for each obstacle introduces a large number of variables to the optimization problem, reducing computational efficiency, especially in cluttered environments. Additionally, direct methods like SQP and IPOPT used by TrajOpt and CBF-Polytopes are relatively slower compared to the DDP-based method AL-iLQR adopted in our framework, as shown in Table II. On the other hand, both Safe-Corridor and our method generate safe trajectories by exploring the relationship between robots and free regions in cluttered environments. However, Safe-Corridor treats the robot as a point and expands obstacles based on the robot’s maximum side length, searching for the reference path using the expanded obstacle information. As demonstrated in Fig. 5, it fails to find the traversable path with the shortest path length, and the trajectory is generated in the shrunk free region (some of the gaps between the free regions and the obstacles are highlighted), leading to a conservative motion. Additionally, it lacks strict safety guarantees and can only provide a safe trajectory when densely sampled waypoints are included in the safety constraints when solving the backend minimum snap problem.
Our method, however, addresses these limitations by ensuring safety through the bi-level solving process with AL-iLQR. We solve the proposed minimum scaling SDP problem along the trajectory of 64 time stamps in parallel after each forward iteration. Each batch takes approximately 6 $\pm$ 3 ms, with the entire AL-iLQR converging in 0.561 s. The results show that our method achieves a shorter trajectory length and faster optimization time in cluttered environments compared to the benchmark methods, while also ensuring safety and non-conservativeness.
@@ V-B Real-World Experiment
In this section, we demonstrate the practicality and robustness of our proposed method by deploying it on a real robot platform Unitree Go1. In the experiment, the quadruped is commanded to traverse through a cluttered indoor area of $2\,\textup{m}\times 6\,\textup{m}$ with several randomly placed obstacles and a hole to simulate the height-constrained situation. We use the onboard Robosense 16-line spinning Lidar to perceive point cloud data for off-line space decomposition and real-time align-based localization. We incorporate additional physical constraints of the quadruped in the trajectory optimization process and implement the kinematic feedforward and feedback controllers for trajectory tracking. The computed body twist commands are subsequently fed into the embedded speed controller to actuate each joint on the robot. The robot, along with the onboard sensor, is approximated as a polytope. The proposed optimization problem (7) is solved in around 15 $\pm$ 3 ms. As depicted in Fig. 6, the quadruped adjusts its pose to adapt to the obstacle configurations, and due to the tightly fitted robot geometry, the quadruped lowers its pose to successfully traverse the height-constrained hole in a non-conservative manner (shown in Fig. 6(g)-6(h)).